1 litre of oxygen and 1 litre of hydrogen are taken in a vessel of 2 litres capacity at NTP. The gases are made to combine by applying electric spark. Assume that water is formed quantitatively. How many gram of water are formed? What is other component present in the vessel and in what weight? If the vessel is now heated to 100°C, what will be the pressure inside the vessel in mm Hg?

Grade:11

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Nikit,

1 lit oxygen and 1 lit hydrogen.

....1/22.4 moles oxygen and 1/22.4 moles hydrogen at NTP..

moles of water formed is 1/22.4 moles...

oxygen left is 1/44.8 moles

weight of oxygen left = 32/44.8 g

weight of water formed is 18/22.4 g

at 373 K,

moles of gases+vapour present = 1/44.8 + 1/22.4 = 3/44.8

P = 3*0.0821*373/(44.8*2) = 1.025 atm

P = 779.254 mm of Hg